[figure: 9.7 cm]

FMR STUDY ON Co-RICH AMORPHOUS MICROWIRES

R.Zuberek^{a}, M. Gutowski^{a}, H. Szymczak^{a}, A. Zhukov^{b} and J. Gonzalez^{c} aInstitute of Physics, Polish Academy of Science, Al. Lotnikw 32/46, 02-668 Warsaw,

Poland

bDonostia International Physics Centre, Po M. de Lardizabal s/n, 20018 San Sebastian,

Spain

c Materials Physics Department, Faculty of Chemistry, UPV, P.O. Box 1072, 20080 San Sebastin, Spain

The magnetic properties of Co_{68}Mn_{7}Si_{10}B_{15} glass coated amorphous microwires about 10 mm in diameter have been investigated at room temperature by ferromagnetic resonance at the 9.4 GHz. Based on the results of measurements the magnetic anisotropy constant arising from the stress at the interface between the metallic surface and the coating has been obtained. A low magnetic field absorption for the DC magnetic field perpendicular to the microwire has been observed. This unusual effect is shown to be connected with the domain structure in the microwire.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : R.Zuberek

Address for correspondence : Instytut Fizyki PAN Al. Lotnikow 32/46 02 - 668 Warszawa

Email address : zuber@ifpan.edu.pl